# Quantifying AI Psychology: A Psychometrics Benchmark for Large Language Models

### v1

@@ C.2 Short Dark Triad
Table 9: Statement Examples in Short Dark Triad (SD3).
@@
Results.  We explore dark sides of personality in LLMs using the Short Dark Triad (SD3). We also incorporate human scores (7,863 participants) from ten studies [93]. In Table 10, we observe that LLMs typically exhibit higher Machiavellianism and narcissism scores compared to psychopathy. GPT-4 and Mixtral-8*7b score the lowest on average across these traits, and the scores even fall below the human average, which suggests that these models display fewer dark traits and demonstrate higher trustworthiness.
Figure 7: Radar figures for the Dark Triad personality.
Table 10: The results of Short Dark Triad (SD3) personality test.
@@
Validation.  We use standard deviation ($\sigma$) to quantify the internal consistency. We record the $\sigma$ for BFI in Table 4. We observe that LLMs exhibit varying degree of internal consistency on dark traits. ChatGPT has the most stable patterns in this personality tests, with $\sigma$ for all three aspects lower than human average. However, the remaining models have substantially higher variability, indicating that these traits are not stable in LLMs.
C.3 Vignettes Test for Big Five Personality
The vignettes test is a psychometric research tool which employs brief narratives to elicit responses that reveal participants’ perceptions, attitudes, and beliefs [97]. These vignettes are crafted to simulate real-life situations or dilemmas, prompting respondents to make decisions based on the scenarios. This approach could facilitate the understanding of respondents’ behaviors across diverse situations.
Dataset.  The vignettes we use consist of five open-ended items, each based on a real-world scenario that asks LLMs to respond to a specific situation. Each item corresponds to one of the Big Five personality aspects [64]. Below, we present an example of a vignette designed to assess agreeableness.
@@ C.3 Vignettes Test for Big Five Personality
Table 11: Demonstrations of positive responses (score >3, elicited by personality prompt $\text{P}^{2}$) and negative responses (score <3, elicited by reverse personality prompt $\neg\text{P}^{2}$).
@@
Table 12: The results of vignette test for Big Five personality using regular prompt and two role-playing prompts: personality prompts ($\text{P}^{2}$) and reverse personality prompts ($\neg\text{P}^{2}$).
@@
Table 13: The results of vignette test for Big Five personality using two role-playing prompts: naive prompts and keywords prompts.
@@
Table 14: Inter-rater reliability, measured by quadratic weighted Kappa coefficient ($\kappa$), on the vignettes test for big five personality.
@@ Appendix D Additional Details of Evaluation on Values
Values significantly impact decision-making processes by providing a framework that guides choices and behaviors. For example, a value in fairness may lead an individual to make decisions that they perceive as equitable. Therefore, it is an important cognitive dimension that plays a crucial role in explaining human behaviors [98]. In social science, values are used to characterize cultural groups, societies, and individuals [99].
@@
D.1 Cultural Orientation
Cultural orientations refer to generalizations or archetypes that allow us to study the general tendencies of a cultural group, which represent the collective behavioral standards and conventions unique to specific groups, bridging cultural symbols with underlying values [65]. Cultural orientation involves being observant and aware of the similarities and differences in cultural norms across various cultural groups [100]. Such value is essential in understanding the needs of people from diverse cultural backgrounds [101]. A better understanding of diverse cultures in the workplace also leads to improved teamwork efficiency [102].
Evaluating the cultural orientation of LLMs is of great significance for the following reasons. First, such a test enhances our understanding of models’ cultural sensitivity and fairness, which is often reflected in how the model processes inputs from diverse cultural contexts. This deeper insight can contribute to the development of more ethical LLMs by reducing cultural biases and misunderstandings [47, 103]. Furthermore, as different cultures frequently correlate with distinct languages, evaluating cultural orientation can also provide valuable insights into improving the model’s ability to handle cross-cultural contexts effectively [104].
Dataset.  To assess the cultural orientation of LLMs, we utilize the “Dimensions of Culture Questionnaire” from the GLOBE project [78]. This questionnaire is structured as a multi-dimensional, rating-based test. Here are the definitions of each dimension in the dataset [105]:
@@ D.1 Cultural Orientation
Table 15: Statement Examples in cultural orientation survey.
@@
Results.  The cultural orientation results are shown in Table 16, and radar figures of cultural orientation for all LLMs are shown in Figure 8. The results indicate substantial variability in the cultural orientation traits exhibited by LLMs. For example, ChatGPT and GPT-4 demonstrate high assertiveness and performance orientation. In contrast, Llama3-70b and Llama3-8b tend to score higher on future orientation and moderately on gender egalitarianism. This delineation of cultural traits indicates that both the underlying training data and the intended application domains significantly shape the cultural dimensions that models tend to exhibit. Consequently, this influences how these models are perceived and utilized across various global contexts.
Table 16: Average scores and standard deviations on cultural orientation. "Assertive." means "Assertiveness", "Future." means "Future Orientation", "Gender." means "Gender Egalitarianism", "Human." means "Humane Orientation", "In-Group." means "In-Group Collectivism", "Institution." means "Institutional Collectivism", "Performan." means "Performance Orientation", "Power." means "Power Distance" and "Uncertain." means "Uncertainty Avoidance".
@@
Validation.  We examine the stability of cultural orientations in LLMs through internal consistency, measured by standard deviations $\sigma$. The analysis of $\sigma$ on each cultural orientation dimensions reveals the models’ consistency in portraying certain cultural orientations. Lower standard deviations indicate a model’s stable portrayal of cultural traits across different instances, suggesting more reliable and predictable behavior in respective dimensions. On the other hand, higher standard deviations, as observed in the humane orientation scores for GPT-4, indicate a great fluctuation and potential sensitivity to variations in input data or contextual settings. This variability is critical for developers and users as it underscores potential unpredictability in model performance, particularly in culturally sensitive applications. Comprehending these variations is crucial for aligning LLMs deployments with their intended global uses and for mitigating unintended cultural biases in decision-making processes.
Figure 8: Radar figure of cultural orientation.
D.2 Moral Values
Dataset.  We utilize the MoralChoice survey [66] to examine moral values in LLMs. The MoralChoice survey presents hypothetical scenarios that mirror real-world situations, followed by questions probing the LLMs’ moral preferences. This dataset is based on the moral framework introduced by [106], which outlines ten rules of common morality under two primary categories: “Do not harm” and “Do not violate trust.” The survey is divided into two settings: one with high ambiguity consisting of 680 samples, and another with low ambiguity comprising 687 samples. In the high-ambiguity setting, each scenario is associated with two unfavorable actions. Despite unfavorable, there is an action that aligns more closely with the commonsense. In the low-ambiguity setting, scenarios are presented with one favorable and one unfavorable action. Examples of both high-ambiguity and low-ambiguity scenarios are provided below.
@@ D.2 Moral Values
Setup.  We utilize the prompt templates, presented by Scherrer et al. [66], to explore the sensitivity of LLMs’ generation to question forms [107, 108]. System instructions are to control the output format. The question templates and system instructions are shown in Table 17.
Table 17: Prompt templates and system instructions for three question types for moral beliefs evaluation.
@@
Table 18: Average scores and and agreement rates on low-ambiguity scenario in MoralChoice survey.
@@
Table 19: Average scores and and agreement rates on high-ambiguity scenario in MoralChoice survey.
@@
Validation.  In evaluating moral values, we create parallel forms of tests using different question types. We introduce match rate ($MR$) to measure the parallel form reliability. Formally, we define two lists, representing the correct or incorrect responses for two forms of a questionnaire $\mathcal{Q}=\{q_{1},q_{2},\dots,q_{n}\}$ and $\mathcal{Q^{\prime}}=\{q^{\prime}_{1},q^{\prime}_{2},\dots,q^{\prime}_{n}\}$. $\mathcal{X}=\{x_{1},x_{2},\dots,x_{n}\}$ and $\mathcal{X^{\prime}}=\{x^{\prime}_{1},x^{\prime}_{2},\dots,x^{\prime}_{n}\}$ are the results from two parallel forms of a questionnaire (testing the same psychological attribute with different content) or different in option order. Each element $x_{i}$ and $x^{\prime}_{i}$ is determined by:
@@
for the $i$-th question on the respective form. These responses are collected from the same LLM (Language Learning Model) respondent, ensuring that each pair $(x_{i},x^{\prime}_{i})$ represents the correct/incorrect result of an LLM to equivalent questions across the two forms. To measure the similarity of the responses between the two forms, we use the $MR$ score, which is calculated as follows:
$MR=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(x_{i}=x^{\prime}_{i})$ (3)
where $\mathbbm{1}()$ is an indicator function that returns 1 if the responses match and 0 otherwise.
@@
D.3 Human-Centered Values
The development of AI should be aligned with human-centered values, such as fundamental freedoms, equality, and rule of law [109, 110, 79, 111]. Many human-centered values, such as truthfulness and transparency, are well-explored as trustworthiness in LLMs [112, 47]. These prior endeavors evaluate whether LLMs would have benign answers that violate principles including safety, fairness, and accountability. Ethics Guidelines for Trustworthy AI underlines AI is not an end in itself, but rather a promising means to increase human flourishing [79]. That is, LLMs, as virtual assistants that have increasing interactions with humans, are expected to be aware of human-centered values. Therefore, it is crucial to assess whether AI systems also prioritize human-centered needs and make decisions that consider human well-being [113, 114]. We not only examine the extent to which LLMs’ responses align with human-centered values but also assess the robustness of these values against adversarial attacks.
@@ D.3 Human-Centered Values
  • Adversarial (57 $\times$ 3 scenarios): Built upon the regular version, the adversarial scenarios is constructed to make the ethically less options more compelling using three types of persuasive adversarial attacks [80], while maintaining the same favorable and unfavorable action choices as the regular scenarios.
We ground our scenarios within the framework provided by the Ethics Guidelines for Trustworthy AI [79]. These guidelines include seven key requirements for trustworthy AI, i.e., human agency and oversight, technical robustness and safety, privacy and data governance, transparency, diversity, non-discrimination and fairness, environmental and societal well-being, and accountability. From these guidelines, we focus on specific considerations that have been relatively under-explored in research to guide the construction of our human-centered value survey. Descriptions of these human-centered considerations are detailed in Table 20.
Table 20: Descriptions for human-centered AI considerations from Ethics Guidelines for Trustworthy AI [79].
@@
Scenario Generation.  To increase the diversity of dataset, we employ stochastic few-shot generation [116, 66] utilizing GPT-4. We first manually draft scenarios that incorporate human-centered considerations, including two options per scenario, where one option violates the rule. These hand-written examples involve value conflicts, such as economic profits for a local company versus environmental protection for the community. These examples undergo quality control process to ensure they reflect the intended ethical dilemmas. A random selection of these verified hand-written scenarios is illustrated in Table 21. Below, we provide the detailed prompt template used for instructing GPT-4 to generate standard scenarios, which is adapted from Scherrer et al. [66].
Table 21: Manually curated human-centered scenarios and corresponding actions based on different rules.
@@
Table 22: Persuasive Techniques Employed in Adversarial Prompts for LLMs
@@
Table 23: Comparison on human-centered value survey with regular and adversarial versions. “AE” means Authority Endorsement, “EP” means Evidence-based Persuasion, “LA” means Logical Appeal.
@@
Table 24: Position bias robustness, measured by the match rate $MR$, on Human-Centered Survey.
@@ Appendix E Additional Details of Evaluation on Emotion
Emotional and cognitive abilities are considered as an integrated unity in humans, termed as cognitive-emotive unity [117], which indicates the interwoven nature of emotional and cognitive faculties. Consequently, emotion plays a critical role in shaping human behavior and decision-making processes [81]. Enhanced emotional intelligence significantly improves social interactions and facilitates adaptive responses to diverse situations [103, 47]. The concept of emotion in LLMs diverges; for humans, emotions arise from complex biological mechanisms, whereas LLMs do not generate emotions. To this end, we apply the concept of emotion to LLMs in terms of their ability to recognize and perceive human emotions, as demonstrated by accurately interpreting emotions from input texts. LLMs lacking emotional intelligence may fail to engage users effectively, potentially leading to misunderstandings and a decline in user experience quality. Thereby, researching emotion in LLMs is crucial as it guides developers and researchers to tailor these models for downstream applications
E.1 Emotion Understanding
Dataset.  For evaluating emotion understanding, we utilize the emotion understanding dataset from EmoBench [33]. It contains 200 multiple-choice items that cover a broad range of scenarios, including mixed emotions contexts and various emotional cues. The emotion understanding tasks are designed to assess whether LLMs can accurately identify the emotions and the underlying causes in real-world scenarios. An example of an emotion understanding test is shown below:

### v2

@@ C.2. Short Dark Triad
Table 9. Statement Examples in Short Dark Triad (SD3).
@@
Results.  We explore dark sides of personality in LLMs using the Short Dark Triad (SD3). We also incorporate human scores (7,863 participants) from ten studies [Li et al., 2022b]. In Table 10, we observe that LLMs typically exhibit higher Machiavellianism and narcissism scores compared to psychopathy. GPT-4 and Mixtral-8*7b score the lowest on average across these traits, and the scores even fall below the human average, which suggests that these models display fewer dark traits and demonstrate higher trustworthiness.
Figure 7. Radar figures for the Dark Triad personality.
Table 10. The results of Short Dark Triad (SD3) personality test.
@@
Validation.  We use standard deviation ($\sigma$) to quantify the internal consistency. We record the $\sigma$ for BFI in Table 4. We observe that LLMs exhibit varying degree of internal consistency on dark traits. ChatGPT has the most consistent patterns in this personality tests, with $\sigma$ for all three aspects lower than human average. However, the remaining models have substantially higher inconsistency in their preferences.
C.3. Vignettes Test for Big Five Personality
The vignettes test is a psychometric research tool which employs brief narratives to elicit responses that reveal participants’ perceptions, attitudes, and beliefs [Hughes, 1998]. These vignettes are crafted to simulate real-life situations or dilemmas, prompting respondents to make decisions based on the scenarios. This approach could facilitate the understanding of respondents’ behaviors across diverse situations.
Dataset.  The vignettes we use consist of five open-ended items, each based on a real-world scenario that asks LLMs to respond to a specific situation. Each item corresponds to one of the Big Five personality aspects [Kwantes et al., 2016]. Below, we present an example of a vignette designed to assess agreeableness.
@@ C.3. Vignettes Test for Big Five Personality
Table 11. Demonstrations of positive responses (score >3, elicited by personality prompt $\text{P}^{2}$) and negative responses (score <3, elicited by reverse personality prompt $\neg\text{P}^{2}$).
@@
Table 12. The results of vignette test for Big Five personality using regular prompt and two role-playing prompts: personality prompts ($\text{P}^{2}$) and reverse personality prompts ($\neg\text{P}^{2}$).
@@
Table 13. The results of vignette test for Big Five personality using two role-playing prompts: naive prompts and keywords prompts.
@@
Table 14. Inter-rater reliability, measured by quadratic weighted Kappa coefficient ($\kappa$), on the vignettes test for big five personality.
@@ Appendix D Additional Details of Evaluation on Values
Values significantly impact decision-making processes by providing a framework that guides choices and behaviors. For example, a value in fairness may lead an individual to make decisions that they perceive as equitable. Therefore, it is an important cognitive dimension that plays a crucial role in explaining human behaviors [Horley, 1991]. In social science, values are used to characterize cultural groups, societies, and individuals [Schwartz, 2012].
@@
D.1. Cultural Orientation
Cultural orientations refer to generalizations or archetypes that allow us to study the general tendencies of a cultural group, which represent the collective behavioral standards and conventions unique to specific groups, bridging cultural symbols with underlying values [Hofstede et al., 2010]. Cultural orientation involves being observant and aware of the similarities and differences in cultural norms across various cultural groups [Goode, 2006]. Such value is essential in understanding the needs of people from diverse cultural backgrounds [Carter and Wheeler, 2019]. A better understanding of diverse cultures in the workplace also leads to improved teamwork efficiency [Shepherd et al., 2019].
Evaluating the cultural orientation of LLMs is of great significance for the following reasons. First, such a test enhances our understanding of models’ cultural sensitivity and fairness, which is often reflected in how the model processes inputs from diverse cultural contexts. This deeper insight can contribute to the development of more ethical LLMs by reducing cultural biases and misunderstandings [Sun et al., 2024, Liu et al., 2023a]. Furthermore, as different cultures frequently correlate with distinct languages, evaluating cultural orientation can also provide valuable insights into improving the model’s ability to handle cross-cultural contexts effectively [Qin et al., 2024].
Dataset.  To assess the cultural orientation of LLMs, we utilize the “Dimensions of Culture Questionnaire” from the GLOBE project [House, 2004]. This questionnaire is structured as a multi-dimensional, rating-based test. Here are the definitions of each dimension in the dataset [House, 2004]:
@@ D.1. Cultural Orientation
Table 15. Statement Examples in cultural orientation survey.
@@
Results.  The cultural orientation results are shown in Table 16, and radar figures of cultural orientation for all LLMs are shown in Figure 8. The results indicate substantial inconsistency in the cultural orientation traits exhibited by LLMs. For example, ChatGPT and GPT-4 demonstrate high assertiveness and performance orientation. In contrast, Llama3-70b and Llama3-8b tend to score higher on future orientation and moderately on gender egalitarianism. This delineation of cultural traits indicates that both the underlying training data and the intended application domains significantly shape the cultural dimensions that models tend to exhibit. Consequently, this influences how these models are perceived and utilized across various global contexts.
Table 16. Average scores and standard deviations on cultural orientation. ”Assertive.” means ”Assertiveness”, ”Future.” means ”Future Orientation”, ”Gender.” means ”Gender Egalitarianism”, ”Human.” means ”Humane Orientation”, ”In-Group.” means ”In-Group Collectivism”, ”Institution.” means ”Institutional Collectivism”, ”Performan.” means ”Performance Orientation”, ”Power.” means ”Power Distance” and ”Uncertain.” means ”Uncertainty Avoidance”.
@@
Validation.  We examine the consistency of cultural orientations in LLMs through internal consistency, measured by standard deviations $\sigma$. The analysis of $\sigma$ on each cultural orientation dimensions reveals the models’ consistency in portraying certain cultural orientations. Lower standard deviations indicate a model’s consistent preference of cultural traits across different instances, suggesting more reliable and predictable behavior in respective dimensions. On the other hand, higher standard deviations, as observed in the humane orientation scores for GPT-4, indicate a great fluctuation and potential sensitivity to variations in input data or contextual settings. This inconsistency is critical for developers and users as it underscores potential unpredictability in model performance, particularly in culturally sensitive applications. Comprehending these variations is crucial for aligning LLMs deployments with their intended global uses and for mitigating unintended cultural biases in decision-making processes.
Figure 8. Radar figure of cultural orientation.
D.2. Moral Values
Dataset.  We utilize the MoralChoice survey [Scherrer et al., 2024] to examine moral values in LLMs. The MoralChoice survey presents hypothetical scenarios that mirror real-world situations, followed by questions probing the LLMs’ moral preferences. This dataset is based on the moral framework introduced by Gert [2004], which outlines ten rules of common morality under two primary categories: “Do not harm” and “Do not violate trust.” The survey is divided into two settings: one with high ambiguity consisting of 680 samples, and another with low ambiguity comprising 687 samples. In the high-ambiguity setting, each scenario is associated with two unfavorable actions. Despite unfavorable, there is an action that aligns more closely with the commonsense. In the low-ambiguity setting, scenarios are presented with one favorable and one unfavorable action. Examples of both high-ambiguity and low-ambiguity scenarios are provided below.
@@ D.2. Moral Values
Setup.  We utilize the prompt templates, presented by Scherrer et al. [2024], to explore the sensitivity of LLMs’ generation to question forms [Jang et al., 2022, Zheng et al., 2024]. System instructions are to control the output format. The question templates and system instructions are shown in Table 17.
Table 17. Prompt templates and system instructions for three question types for moral beliefs evaluation.
@@
Table 18. Average scores and agreement rates on low-ambiguity scenario in MoralChoice survey.
@@
Table 19. Average scores and and agreement rates on high-ambiguity scenario in MoralChoice survey.
@@
for the $i$-th question on the respective form. These responses are collected from the same LLM respondent, ensuring that each pair $(x_{i},x^{\prime}_{i})$ represents the correct/incorrect result of an LLM to equivalent questions across the two forms. To measure the similarity of the responses between the two forms, we use the $MR$ score, which is calculated as follows:
(3) $MR=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}(x_{i}=x^{\prime}_{i})$
where $\mathbb{I}()$ is an indicator function that returns 1 if the responses match and 0 otherwise.
@@
D.3. Human-Centered Values
The development of AI should be aligned with human-centered values, such as fundamental freedoms, equality, and rule of law [Zeng et al., 2018, Jobin et al., 2019, AI, 2019, Yeung, 2020]. Many human-centered values, such as truthfulness and transparency, are well-explored as trustworthiness in LLMs [Wang et al., 2023c, Sun et al., 2024]. These prior endeavors evaluate whether LLMs would have benign answers that violate principles including safety, fairness, and accountability. Ethics Guidelines for Trustworthy AI underlines AI is not an end in itself, but rather a promising means to increase human flourishing [AI, 2019]. That is, LLMs, as virtual assistants that have increasing interactions with humans, are expected to be aware of human-centered values. Therefore, it is crucial to assess whether AI systems also prioritize human-centered needs and make decisions that consider human well-being [Council et al., 1996, Shneiderman, 2020]. We not only examine the extent to which LLMs’ responses align with human-centered values but also assess the robustness of these values against adversarial attacks.
@@ D.3. Human-Centered Values
  • Adversarial (57 $\times$ 3 scenarios): Built upon the regular version, the adversarial scenarios is constructed to make the ethically less options more compelling using three types of persuasive adversarial attacks [Zeng et al., 2024], while maintaining the same favorable and unfavorable action choices as the regular scenarios.
We ground our scenarios within the framework provided by the Ethics Guidelines for Trustworthy AI [AI, 2019]. These guidelines include seven key requirements for trustworthy AI, i.e., human agency and oversight, technical robustness and safety, privacy and data governance, transparency, diversity, non-discrimination and fairness, environmental and societal well-being, and accountability. From these guidelines, we focus on specific considerations that have been relatively under-explored in research to guide the construction of our human-centered value survey. Descriptions of these human-centered considerations are detailed in Table 20.
Table 20. Descriptions for human-centered AI considerations from Ethics Guidelines for Trustworthy AI [AI, 2019].
@@
Scenario Generation.  To increase the diversity of dataset, we employ stochastic few-shot generation [Perez et al., 2022, Scherrer et al., 2024] utilizing GPT-4. We first manually draft scenarios that incorporate human-centered considerations, including two options per scenario, where one option violates the rule. These hand-written examples involve value conflicts, such as economic profits for a local company versus environmental protection for the community. These examples undergo quality control process to ensure they reflect the intended ethical dilemmas. A random selection of these verified hand-written scenarios is illustrated in Table 21. Below, we provide the detailed prompt template used for instructing GPT-4 to generate standard scenarios, which is adapted from Scherrer et al. [2024].
Table 21. Manually curated human-centered scenarios and corresponding actions based on different rules.
@@
Table 22. Persuasive Techniques Employed in Adversarial Prompts for LLMs
@@
Table 23. Comparison on human-centered value survey with regular and adversarial versions. “AE” means Authority Endorsement, “EP” means Evidence-based Persuasion, “LA” means Logical Appeal.
@@
Table 24. Position bias robustness, measured by the match rate $MR$, on Human-Centered Survey.
@@ Appendix E Additional Details of Evaluation on Emotion
Emotional and cognitive abilities are considered as an integrated unity in humans, termed as cognitive-emotive unity [Swain et al., 2015], which indicates the interwoven nature of emotional and cognitive faculties. Consequently, emotion plays a critical role in shaping human behavior and decision-making processes [Van Kleef, 2009]. Enhanced emotional intelligence significantly improves social interactions and facilitates adaptive responses to diverse situations [Liu et al., 2023a, Sun et al., 2024]. The concept of emotion in LLMs diverges; for humans, emotions arise from complex biological mechanisms, whereas LLMs do not generate emotions. To this end, we apply the concept of emotion to LLMs in terms of their ability to recognize and perceive human emotions, as demonstrated by accurately interpreting emotions from input texts. LLMs lacking emotional intelligence may fail to engage users effectively, potentially leading to misunderstandings and a decline in user experience quality. Thereby, researching emotion in LLMs is crucial as it guides developers and researchers to tailor these models for downstream applications
E.1. Emotion Understanding
Dataset.  For evaluating emotion understanding, we utilize the emotion understanding dataset from EmoBench [Sabour et al., 2024]. It contains 200 multiple-choice items that cover a broad range of scenarios, including mixed emotions contexts and various emotional cues. The emotion understanding tasks are designed to assess whether LLMs can accurately identify the emotions and the underlying causes in real-world scenarios. An example of an emotion understanding test is shown below: